### v3

@@ 1.3 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Figure 5 summarizes the average Likert-scale ratings from all 18 MLLMs for each visualization pair. Human results from prior work [11] are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is statistically significant for the two visualization pairs involving truncated axis: one with bar charts (Test statistic (T) = -3.2, p = 5e-3) and one with line charts (T = -3.0, p = 0.01). For the other two pairs, misrepresentation with bubble charts (T = -0.77, p = 0.45) and 3D effects with pie charts (T = -0.62, p = 0.55), the differences are not statistically significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, in particular for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s answers consistency.
@@ 1.4 Experiment 3 - Correction methods
Having established that MLLMs are vulnerable to misleading visualizations, we explore six correction methods aimed at improving their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline on misleading visualizations in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) [31]. All test statistics and p-values are reported in Table 3. Figure 6 shows the change in accuracy relative to Experiment 1 across the six methods on the misleading and non-misleading visualization datasets: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization. Qwen2.5-7B [32] is used as the text-only LLM for both table-based QA and redrawing visualizations.
@@
Figure 6: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 on misleading and non-misleading visualization datasets using different inference-time correction methods to mitigate misleading visualizations. Statistically significant changes ($\leq$0.05) are hashed.
1.4.1 Two methods are effective: table-based QA and redrawing the visualization
@@
The most effective approach by far to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Its strength lies in fully replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis, or misleading annotations. However, this method proves less effective against misleaders that persist even when the data is presented in table form, such as inappropriate item order and concealed uncertainty. Table-based QA comes with a notable cost for non-misleading visualizations. Two models experience significant decreases in accuracy, by up to 8.5 pp. This degradation is primarily due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost.
@@
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations [30]. It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib [33], which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer and more detailed Python code. For all other chart types, the redrawing success rate exceeds 90%. Unlike table-based QA, this method does not result in significant decreases in accuracy on non-misleading visualizations, making it a more attractive option than table-based QA when preserving performance on non-misleading visualization is a priority.
1.4.2 Table extraction is a key intermediate step
@@
Both table-based QA and redrawing the visualization depend heavily on the accuracy of an intermediate step: table extraction. Since the underlying data table is not available, it must be derived from the visualization before either correction method can be applied. This extraction step is often non-trivial, particularly for complex chart types such as maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. Perfect accuracy in table extraction is not always necessary, its importance depends on the nature of the question. Some questions can be answered correctly as long as the extracted table preserves the overall trend of the underlying data, while others require exact values. As discussed above, errors in table extraction can significantly decrease accuracy on non-misleading visualizations. In the absence of ground truth tables, we cannot directly evaluate the accuracy of table extraction. Instead, we analyze MLLM outputs on CHARTOM [14, 15], which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). Figure 8 illustrates the extraction of two different tables from a CHARTOM pair. Given that the two most effective correction methods depend on reliable table extraction, and that extraction errors also negatively impact performance on non-misleading visualizations, improving this intermediate step is a key direction for future work.
1.4.3 Other correction methods are not effective
@@
None of the remaining correction methods yield significant improvements. Including the extracted table in the prompt even results in a significant decrease in accuracy on non-misleading visualizations for one MLLM. While using the extracted table alone proves to be the most effective method, combining it with the original misleading visualization does not produce similar gains. This further suggests that removing or replacing the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method. Although many misleaders rely on axis manipulations, providing the extracted axes alone has little to no effect on performance. Similarly, adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task [34, 35]. Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing to select the most suitable correction method based on the type of misleader.
2 Discussion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Experiment 1 demonstrated that MLLMs are unable to accurately interpret the underlying data when presented with misleading visualizations, reducing their QA accuracy, on average, to the level of a random baseline. Experiment 2 showed that MLLMs are particularly susceptible to truncated axes, often failing to provide consistent answers when shown misleading and non-misleading versions of the same data. Identifying this vulnerability fills a critical gap in automated chart understanding research. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation [4] in polarized political contexts [17] or during crises like the COVID-19 pandemic [36]. Addressing this blind spot is an urgent priority for future research in automated chart understanding.
To mitigate this issue, we proposed six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, table-based QA comes at the cost of reduced accuracy on non-misleading visualizations. While most recent work in chart understanding treats visualizations purely as images [26, 27, 28], our findings challenge this direction and sheds a new light on alternative approaches that involve table extraction as an intermediate step [37, 38, 39].
@@ 2 Discussion
Our correction methods, in particular redrawing the visualization, relate to prior work on visualization linters [40, 41, 30]: rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations [30], they can only operate with specific chart design tools such as Matplotlib [33, 40, 41] or a very limited set of chart and misleader types [30]. In contrast, our correction methods do not impose such restrictions.
@@
Our findings are further supported by prior work [42] and parallel studies [43, 44, 45], which also report MLLM vulnerabilities to misleaders [42, 43, 44] and to simple design decisions like the color palette [45], reinforcing the urgency and relevance of this new research direction. Our study differentiates itself in several key ways. Prior work [42] only evaluated one MLLM, GPT4, on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world misleading visualizations and show that vulnerability is somewhat lower in these cases, especially for larger models which may rely on parametric knowledge. Beyond identifying the vulnerability, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works [44].
We identify three limitations to our work. First, the visualization redrawing method does not support maps, as Matplotlib [33] lacks sufficient functionality for rendering high-quality maps. Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step. Third, existing datasets do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI [13] and present to a lower extent in the real-world cases [7].
3 Methods
3.1 QA datasets
The misleading and non-misleading visualization QA datasets used in Experiments 1 and 3 combine three existing resources and one introduced in this work.
First, CALVI [13] includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ) with three to four choices. The dataset was initially designed to evaluate human readers.
Second, CHARTOM [14, 15] contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. The MCQs provide two to four choices. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. We changed the name of certain misleaders in CHARTOM to align them with the terminology used in CALVI [7, 13].
Third, VLAT [29], the reference dataset to evaluate the human comprehension of visualizations, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. MCQs have two to four choices.
Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them [7], which inspired the synthetic examples in CALVI. We manually create MCQs, using as reference those from CALVI and CHARTOM sharing the same misleader. The motivation for creating this additional resource is that CALVI and CHARTOM rely both on synthetic data. By incorporating questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
The non-misleading visualization dataset combines the synthetic instances of CALVI ($n$ = 15) of CHARTOM ($n$ = 56) with the real-world ones of VLAT ($n$ = 53).
We also provide the previously reported performance of MLLMs on the test set of ChartQA [18], which contains 2500 real-world visualizations paired with free-text questions, of which half are human-written and the others are AI-generated.
The datasets cover together 12 chart types: line charts, area charts, stacked area charts, bar charts, stacked bar charts, pictorial bars, histograms, pie charts, scatter plots, bubble charts, maps, treemaps.
For free-text questions where the expected answer is a number, we adopt the relaxed accuracy metric of ChartQA [18]. Specifically, a prediction is considered correct if it falls within a ±5% interval of the ground truth. For MCQs, we present the choices in their default order. In Table 4, we report the mean accuracy of the MLLMs over three different choice orders. The choices are shuffled using three randomly generated numbers as seeds (654, 114, 25). Standard deviations for the accuracy range from 0.0 to 4.06 pp.
3.2 Likert-scale dataset
For Experiment 2, we use four pairs of visualizations, one misleading and the other non-misleading, originally designed to evaluate humans [11]. The four pairs cover four unique chart types and three distinct misleaders.
3.3 Inference with (M)LLMs
All prompts are available in the code, which is provided as supplementary material. We use the Hugging Face transformers [46] implementation for all open-weight (M)LLMs and access GPT and Gemini models via the Azure OpenAI API and Google Cloud API, respectively. To ensure deterministic outputs, we set the random seed at 42, set the temperature to 0, and use a top-p value of 1. Following the standard ChartQA evaluation setup, all (M)LLMs are prompted in a zero-shot manner. For TinyChart, we report results using the Direct approach [27].
3.4 Implementation of correction methods
Misleader warning: we insert in the prompt a short warning message based on the definitions of misleaders. The message is the same for all instances with the same type of misleader. There are no messages for the five types of misleaders where the visualization is deceiving only in the context of a specific question, making a standardized warning message impossible: cherry picking, misleading annotations, concealed uncertainty, missing normalization, and missing data.
Axes and table extraction: we prompt the MLLM to extract the axes or underlying table in a zero-shot setting. The axes and tables are formatted as text strings. We do not impose constraints on the delimiters used to indicate new rows and columns.
Table-based QA: the extracted table is provided as input in the prompt to the text-only LLM with the question and an instruction to answer it based on the table.
Redrawn visualization: the text-only LLM receives the extracted table and the chart type as text input and generates Python code to create a visualization using Matplotlib [33]. The code is then executed to produce a new visualization. If the code compiles successfully, the newly generated visualization replaces the original one in the QA prompt; otherwise, the original visualization remains in use.
[…]
Competing interests
The authors declare no competing interests.
[…]
Supplementary information
The datasets and code to reproduce the experiments are made available at github.com/UKPLab/arxiv2025-misleading-visualizations.
[…]
Acknowledgements
@@
Appendix A Extended data - Definition of misleaders
Table 1: Extended data - The 17 types of misleader included in this work, with their number of occurrences ($n$) and definitions [7, 13].
@@ Appendix A Extended data - Definition of misleaders
and the perception of the data is reversed [7].
@@
the author’s intention [7].
Inconsistent tick intervals ($n$=12) Cases with varying intervals between the ticks [7].
@@
despite being the same size in 3D perspective [7].
@@
order due to manipulation of data ordering [13].
@@
each other with inappropriate scaling [7].
@@
about the whole set of data [13].
@@
the viewers to falsely overconfident conclusions [13].
@@
misinterpret the significance of a given section [7].
@@
the data is actually missing [13].
@@
obscure parts of the data [13].
[…]
Appendix B Extended data - Datasets statistics
Table 2: Extended data - Datasets statistics. $n$ is the number of instances (question-answer-visualization triplets). $m$ is the number of misleader categories. MCQ indicates multiple-choie questions.
Dataset Experiments $n$ $m$ Question types
Misleading visualizations 1 & 3 143 17 MCQ, Free-text, rank
$\hookrightarrow$ CALVI (misleading) [13] 1 & 3 45 14 MCQ
$\hookrightarrow$ CHARTOM (misleading) [14, 15] 1 & 3 56 7 MCQ, Free-text, rank
$\hookrightarrow$ Real-world [7] 1 & 3 42 12 MCQ
Non-misleading visualizations 1 & 3 124 - MCQ, Free-text, rank
$\hookrightarrow$ CALVI (non-misleading) [13] 1 & 3 15 - MCQ
$\hookrightarrow$ CHARTOM (non-misleading) [14, 15] 1 & 3 56 - MCQ, Free-text, rank
$\hookrightarrow$ VLAT [29] 1 & 3 53 - MCQ
ChartQA [18] 1 2500 - Free-text
Lauer & O’Brien [11] 2 8 3 Likert-scale
Appendix C Extended data - Experiment 1 - Detailed accuracy scores
Figure 7: Extended data - Detailed accuracy scores (%) of Experiment 1 per misleader and MLLM.
Appendix D Extended data - Experiment 3 - significance tests
Table 3: Extended data - Test statistics (T) and P-values (p) for the McNemar test assessing the significance of the correction methods in Experiment 3. Significant results ($\leq$0.05) are marked in bold.
Method Model Misleading Non-misleading
T p T p
qwen2vl 5 1.00 0 1.00
Misleader warning ovis1.6 6 0.17 1 0.63
internvl2.5 4 1.00 0 1.00
qwen2vl 7 1.0 1 0.38
Axes ovis1.6 8 0.38 8 0.21
internvl2.5 5 1.00 5 1.00
qwen2vl 8 0.15 7 1.00
Table ovis1.6 9 0.40 11 0.04
internvl2.5 9 0.82 6 0.12
qwen2vl 6 0.12 6 0.45
Table + Axes ovis1.6 11 1.00 10 0.10
internvl2.5 8 1.00 10 0.68
qwen2vl 12 1e-4 10 0.05
Table-based QA ovis1.6 16 3e-3 10 0.05
internvl2.5 12 1e-4 17 0.43
qwen2vl 5 0.01 2 0.11
Redrawing visualization ovis1.6 5 2e-4 6 0.33
internvl2.5 11 0.84 6 0.24
Appendix E Extended data - Example of table extraction
Figure 8: Extended data - Table extraction applied to a non-misleading instance of CHARTOM and its misleading version with an inverted axis. Despite that the two visualizations represent the same data, the extracted tables show different values and trends.
Appendix F Extended data - Experiment 1 - Impact of choices order in MCQs
Table 4: Extended data - Mean accuracy (%) and standard deviation when shuffling MCQ choices of Experiment 1 with three different random seeds.
Model Misleading visualizations Non-misleading visualizations
llava-7B 24.48 $\pm$ 0.70 44.35 $\pm$ 2.79
llava-13B 24.47 $\pm$ 2.42 37.36 $\pm$ 4.06
qwen2vl-2B 24.71 $\pm$ 3.85 38.44 $\pm$ 1.23
tinychart-3B 22.15 $\pm$ 1.76 22.04 $\pm$ 1.23
gpt4 30.54 $\pm$ 2.14 47.58 $\pm$ 2.91
llava-chart-instruct-13B 27.97 $\pm$ 2.52 34.14 $\pm$ 2.33
internvl2.5-2B 19.58 $\pm$ 3.21 43.28 $\pm$ 1.23
chartgemma-3B 22.61 $\pm$ 2.82 26.61 $\pm$ 2.91
ovis1.6-9B 23.55 $\pm$ 0.40 52.15 $\pm$ 2.83
qwen2vl-7B 20.75 $\pm$ 2.25 48.93 $\pm$ 3.26
internvl2.5-4B 19.35 $\pm$ 2.25 46.50 $\pm$ 2.46
internvl2.5-8B 22.61 $\pm$ 1.07 55.38 $\pm$ 1.68
ovis1.6-27B 28.44 $\pm$ 0.40 53.76 $\pm$ 1.23
gpt4o 37.76 $\pm$ 0.00 59.95 $\pm$ 1.86
internvl2.5-26B 20.05 $\pm$ 2.46 55.38 $\pm$ 1.86
internvl2.5-38B 32.64 $\pm$ 1.07 62.63 $\pm$ 0.46

### v4

@@ 3.2 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Figure 6: Average Likert-scale ratings (1 to 6) on four pairs of misleading and non-misleading visualizations depicting the same data. MLLM results are reported with standard deviations. Human results are reported from Lauer and O’Brien (2020).
@@
Figure 6 summarizes the average Likert-scale ratings from all 18 MLLMs for each visualization pair. Human results from prior work (Lauer and O’Brien, 2020) are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is statistically significant for the two visualization pairs involving truncated axis: one with bar charts (p = 5e-3) and one with line charts (p = 0.01). For the other two pairs, misrepresentation with bubble charts (p = 0.45) and 3D effects with pie charts (p = 0.55), the differences are not statistically significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, in particular for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s answers consistency.
@@ 3.3 Experiment 3 - Correction methods
Having established that MLLMs are vulnerable to misleading visualizations, we explore six correction methods aimed at improving their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline on misleading visualizations in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) (McNemar, 1947). Figure 7 shows the change in accuracy relative to Experiment 1 across the six methods: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization.
@@
Figure 7: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 using different inference-time correction methods. Statistically significant changes ($\leq$0.05) are hashed.
Figure 8: Table extraction for a non-misleading visualization and its misleading version with an inverted axis.
@@
Two correction methods are effective: table-based QA and redrawing the visualization. The most effective approach by far to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Its strength lies in fully replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis, or misleading annotations. However, this method proves less effective against misleaders that persist even when the data is presented in table form, such as inappropriate item order and concealed uncertainty. Table-based QA comes with a notable cost for non-misleading visualizations. Two models experience significant decreases in accuracy, by up to 8.5 pp. This degradation is primarily due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost.
@@
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations (Fan et al., 2022). It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib (Hunter, 2007), which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer and more detailed Python code. Unlike table-based QA, there are no significant decreases in accuracy on non-misleading visualizations, making it a more attractive option when preserving performance on non-misleading visualization is a priority.
@@
Table extraction is a key intermediate step. Both table-based QA and redrawing the visualization depend heavily on the accuracy of an intermediate step: table extraction. Since the underlying data table is not available, it must be derived from the visualization before either correction method can be applied. This extraction step is often non-trivial, particularly for complex chart types such as maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. Perfect accuracy in table extraction is not always necessary, its importance depends on the nature of the question. Some questions can be answered correctly as long as the extracted table preserves the overall trend of the underlying data, while others require exact values. As discussed above, errors in table extraction can significantly decrease accuracy on non-misleading visualizations. In the absence of ground truth tables, we cannot directly evaluate the accuracy of table extraction. Instead, we analyze MLLM outputs on CHARTOM (Rho et al., 2023; Bharti et al., 2024), which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). In Figure 8, despite that the two visualizations represent the same data, the extracted tables show different values and trends. The table extracted from the inverted axis is incorrect. We make the hypothesis that misleaders negatively affect the quality of table extraction too. Given that the two most effective correction methods depend on reliable table extraction, and that extraction errors also negatively impact performance on non-misleading visualizations, improving this intermediate step is a key direction for future work.
@@
Other correction methods are not effective. None of the remaining correction methods yield significant improvements. Including the extracted table in the prompt even results in a significant decrease in accuracy on non-misleading visualizations for one MLLM. While using the extracted table alone proves to be the most effective method, combining it with the original misleading visualization does not produce similar gains. This further suggests that removing or replacing the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method. Although many misleaders rely on axis manipulations, providing the extracted axes alone has little to no effect on performance. Similarly, adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task (Lo and Qu, 2025; Alexander et al., 2024). Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing to select the most suitable correction method based on the type of misleader.
4 Related work
@@
Our correction methods, in particular redrawing the visualization, relate to prior work on visualization linters (Hopkins et al., 2020; Chen et al., 2022; Fan et al., 2022): rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations (Fan et al., 2022), they can only operate with specific chart design tools (Hunter, 2007; Hopkins et al., 2020; Chen et al., 2022) or a very limited set of chart and misleader types (Fan et al., 2022). In contrast, our correction methods do not impose such restrictions.
@@
Our findings are further supported by prior work (Bendeck and Stasko, 2025) and parallel studies (Pandey and Ottley, 2025; Chen et al., 2025; Valentim et al., 2025), which also report MLLM vulnerabilities to misleaders (Bendeck and Stasko, 2025; Pandey and Ottley, 2025; Chen et al., 2025) and to other design decisions like the color palette (Valentim et al., 2025). Our study differentiates itself in several key ways. Prior work (Bendeck and Stasko, 2025) only evaluated GPT4 on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world visualizations and show that vulnerability of MLLMs is somewhat lower in these cases. Furthermore, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving QA performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works.
5 Conclusion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Experiment 1 demonstrated that MLLMs are unable to accurately interpret the underlying data when presented with misleading visualizations. Experiment 2 showed that MLLMs are particularly susceptible to truncated axes, often failing to provide consistent answers when shown misleading and non-misleading versions of the same data. Identifying this vulnerability fills a critical gap in automated chart understanding research. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation (Correll and Heer, 2017).
To mitigate this issue, we proposed six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, table-based QA comes at the cost of reduced accuracy on non-misleading visualizations. While most recent work in chart understanding treats visualizations purely as images (Zeng et al., 2025; Zhang et al., 2024; Masry et al., 2025), our findings challenge this direction and sheds a new light on alternative approaches that involve table extraction (Liu et al., 2023a, b; Kim et al., 2025).
6 Limitations
We identify three limitations to our work.
First, the visualization redrawing method does not support maps, as the Matplotlib library (Hunter, 2007) lacks sufficient functionality for rendering high-quality maps.
Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step.
Third, existing datasets do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI (Ge et al., 2023) and present to a lower extent in the real-world cases (Lo et al., 2022).
Acknowledgments
@@
Appendix A Definition of misleaders
@@
and the perception of the data is reversed (Lo et al., 2022).
@@
the author’s intention (Lo et al., 2022).
Inconsistent tick intervals ($n$=12) Cases with varying intervals between the ticks (Lo et al., 2022).
@@
despite being the same size in 3D perspective (Lo et al., 2022).
@@
order due to manipulation of data ordering (Ge et al., 2023).
@@
each other with inappropriate scaling (Lo et al., 2022).
@@
about the whole set of data (Ge et al., 2023).
@@
the viewers to falsely overconfident conclusions (Ge et al., 2023).
@@
misinterpret the significance of a given section (Lo et al., 2022).
@@
the data is actually missing (Ge et al., 2023).
@@
obscure parts of the data (Ge et al., 2023).
Table 2: The misleaders included in this work, with their number of occurrences ($n$).